OpenStudy (anonymous):

Dana looks around at an assembly. She notices her younger sister to her right and her older brother 21 feet ahead of her. If Dana's brother and sister are 35 feet apart, how far apart are Dana and her younger sister?

OpenStudy (anonymous):

 D = Dana, S = Sister, B = Brother This is a right triangle, so we can use Pythagorean's Theorem, which states that \[a^2 + b^2 = c^2\] Where "a" and "b" are the legs of the right triangle and "c" is the hypotenuse. Now, substitute in the values that you already know. \[21^2 + b^2 = 35^2\] Substitute \[441 + b^2 = 1225\] Simplify \[b^2 = 784\] Subtract 441 from both sides \[b = 28\] Take the square root of both sides Dana is 28 feet away from younger sister. Hope I helped, and please let me know if you have any more questions :D